OpenStudy (anonymous):

.A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (johnweldon1993):

When you have an absolute value bar like that...you really have 2 equations... x - 0.525 = 0.01 + 0.525 + 0.525 ------------------- x = 0.535 and x - 0.525 = -0.01 +0.525 +0.525 ------------------- x = 0.515 Make sense? So altogether it would be between those numbers... \[\large 0.515\le x \le 0.535\]